OpenStudy (anonymous):

Tell whether the sequence is arithmetic . If it is , what is the common difference ? -19,-11,-3, 5 , yes ; 5 yes ; 6 yes ; 8 no

OpenStudy (anonymous):

I'm going to go with No.

OpenStudy (anonymous):

@Mateaus

OpenStudy (anonymous):

This is the hardest subject I every had D: mhmm

OpenStudy (anonymous):

I see a pattern... it's increasing by a number.

OpenStudy (uri):

@crystal201 what's so hard about this? did you try solving it?

OpenStudy (anonymous):

or it could be 8

OpenStudy (uri):

it is

OpenStudy (anonymous):

so if it's increasing constantly by the same number we can conclude that the answer is... :3

OpenStudy (anonymous):

yes i didn't realize that it was 8 till now . cause i wasn't looking at the numbers increasing .

OpenStudy (anonymous):

so the correct answer is yes;8 !

OpenStudy (anonymous):

mhm

OpenStudy (anonymous):

how hard can it be? are you adding the same number each time? and if so, what is that number?

OpenStudy (anonymous):

11-19 = 8 11-3=8

OpenStudy (anonymous):

-3+5=8

OpenStudy (anonymous):

am I correct? lol

OpenStudy (anonymous):

if it is not obvious from your eyes, then do what @crystal201 wrote above subtract the second term from the first twice yes you are right

OpenStudy (anonymous):

well no your answer is right, your method is wrong

OpenStudy (anonymous):

thanks so much :p @satellite73

OpenStudy (anonymous):

\[\huge11-(-19)=8\\ -3-(-11)=8\]

OpenStudy (uri):

this girl is on fire.... she's walkin on fireeeeeeeee.

OpenStudy (anonymous):

thanks everyone :p